OpenStudy (anonymous):

Write these products as sums: cos(2x)sin(5x)

OpenStudy (anonymous):

@AravindG @Mertsj @amistre64

OpenStudy (john_es):

\[\cos A\sin B=\frac{1}{2}(\sin(A+B)-\sin(A-B))\Rightarrow\\ \Rightarrow \cos 2x\sin 5x=\frac{1}{2}(\sin(7x)-\sin(-3x))=\\ =\frac{1}{2}(\sin(7x)+\sin(3x))\]